Personal Details

Name:  Conall O’Brien  Email:  conall@conall.net
Address: 21 Meadow Close,  Phone (PSTN):  01 6535148
Blackrock,  Phone (VoIP):  sip:31313@blueface.ie
Co Dublin,  Personal Website: http://www.conall.net

Education

October 2001 - June 2006  Computer Science BA(Mod)  Dublin University, Trinity College
September 1996 - June 2001 Secondary Education  Blackrock College

Awards

1997: Blackrock College Computer Studies Award
1999: Blackrock College Construction Studies Award

Qualifications

1999:  European Computer Driver’s License (ECDL)
2001:  Leaving Certificate
2006:  BA (Mod) Hons

Affiliations

Committee Member  System Administrators Guild of Ireland (SAGE-IE)
Member  Irish Network Operaters Group (IE-nog)
Member  Irish Linux User Group (ILUG)
Student Member  Irish Computer Society (ICS)
Student Member  USENIX, the Advanced Computing Systems Association
Student Member  System Administrators Guild (SAGE)
Student Member  League of Professional System Administrators (LOPSA)
Package Maintainer  Debian Project

Objective

I intend to continue to increase and expand my computer knowledge and skills, in addition to using my problem solving abilities to full effect. I also intend to obtain IT networking qualifications, starting with a Cisco Certified Network Associate (CCNA) certification and to actively contribute to the Debian GNU/Linux, Ubuntu Linux and FreeBSD Open Source projects.

Professional Experience

March 1998 -  InfoCAD Data LTD  Systems and Network Administrator

June 2006

Responsibilities: Administration of Office and Client IT Infrastructure; User and Client

Support; IT Equipment Purchasing; IT Security Duties.

March 1998 -  Room 205 Computer Room Systems Administrator
August 2002:  Ballinteer Community School
Responsibilities: Administration of IT Systems for Special Needs Resource Centre;

IT Equipment Purchasing.

July 2004 -  IT Dept  Systems and Network Engineer
August 2004:  Critical Path  (Temporary)
Responsibilities: Administration of IT Infrastructure; Migration of Critical Servers;

500+ User Support.

July 2004 -  School of Mathematics  Systems and Network Administrator
June 2006:  Dublin University

Trinity College

Responsibilities: Administration of IT Infrastructure; Liaising with IT Staff

from other College Depts; User Support for 2000+ Dept Staff and Students; IT Security Duties.

June 2006 -  Google Ireland  Systems Administrator
Current  Gordon House

Barrow St Dublin 4

Responsibilities: Administration of internal systems, Engineer interviewing

On call and user support duties.

References

References also available upon request.

Interests

Architecture; Open Source Software; Innovative Technology; Mobile Communications; Networking; Voice over IP (VoIP); Computer Security; Multimedia; Furniture Making.

Computer Experience

Operating Systems

Familiar with administration of Microsoft Windows (3.1x, 9x, ME, XP, NT, 2000, XP, 2003), MS-DOS O/S, GNU/Linux Distributions (Debian, Ubuntu, Red Hat/Fedora,Ensim, Mandrake), Mac OS X, various flavours of Unix (FreeBSD, NetBSD, Solaris, Solaris x86, HP-UX), Palm OS 5 and Cisco IOS.

Hardware

Experience working with ATA, Serial ATA and SCSI based disk systems, copper and fibre optic networking, and Fibre Channel connected SANs. Familiar with x86, AMD64/EM64T, PowerPC, Sparc and Alpha architectures. Experience working with Cisco switches, Cisco 25xx access routers and Cisco 79xx series IP phones.

Networking

Experience using Apache, BIND, Squid, Samba, NFS, Postfix, Dovecot, Asterisk PBX, MySQL, Active Directory, Firewalls, 802.11 a/b/g wireless networking, routing and subnetting, IPv4 and IPv6 unicast and multicast networking.

Programming and Markup Languages

Experience with Java (J2EE, WFC, Swing) PHP, Motorola 68000 Assembly, QBASIC, C, C++ (STL, MFC, Qt), JavaScript, VHDL, Eiffel (SmartEiffel and SmallEiffel), Python, Perl and Shell Scripting (Bash).

Able to write validating HTML (up to 4.01) and XHTML (up to 1.1 Strict) - conforming to accessibility guidelines (WCAG), CSS (up to CSS 2), XML, LaTeX, SQL (MySQL, Oracle, PostgreSQL) and Web Services (SOAP and XMLRPC).

Legal

This CV is copyright Conall O’Brien, 2005 - 2006. You may not modify this CV. You may not distribute this CV to other parties without my express, written permission. In particular, I do not wish to be represented by any agency or third party and consent only to the use and distribution of this CV for the purposes of the immediate position(s) for which I have applied. If you have received this CV from an agent purporting to represent me, please contact me at the above email address or phone number.